Let X be an exponential random variable such that P(X < 27) = P(X > 27). Calculate E[X|X > 23].
Answer #1

let xnenpla) 21 > facere de *** Oiw. prxs27] = p*>27] da 27 deraan du = stenta An 27 an 27 C ten 10 % ? terde br -> -X - 271= 238238 - 58o-25*2=63+) €254 = (23+ 22 23. 2 23 luz 2/27 27 27 ) = (23+ en 2